Olsen Outfitters Inc. believes that its optimal capital structure consists of 70% common equity and 30% debt, and its tax rate is 40%. Olsen must raise additional capital to fund its upcoming expansion. The firm will have $1 million of retained earnings with a cost of rs = 15%. New common stock in an amount up to $7 million would have a cost of re = 18%. Furthermore, Olsen can raise up to $4 million of debt at an interest rate of rd = 9% and an additional $3 million of debt at rd = 12%. The CFO estimates that a proposed expansion would require an investment of $8.7 million. What is the WACC for the last dollar raised to complete the expansion? Round your answer to two decimal places.
Answer #1

require investment = $8.7 million

equity portion = 70%

= 8.7 x 0.7 = $6.09 million

Debt portion = 30%

= 8.7 x 0.3 = $2.61 million

Cost of Debt:

cost of Debt up to $4 million = 9%

tax rate = 40%

after tax cost of debt = r(1-t)

where r = rate , t = tax rate

so after tax cost of debt = 9(1 - 0.4) = 5.4%

Debt portion = 30% or 0.3

Cost of Equity:

cost of retained earnings = 15%

cost of raising new common stock = 18%

so from above it is cheaper to utilize retained earnings first and then for the remaining balance new stock should be issued

Retained earnings = $1 million

weight of retained earnings = 1 / 8.7 = 0.1149

balance to be raised as common stock = 6.09 - 1 = $5.09 million

weight of common stock = 5.09 / 8.7 = 0.5851

WACC:

Weighted average cost of capital = Wd*Rd + Ws*Rs + We*Re

Where, Wd , Ws , We = weights of Debt , Retained earnings , New common stock

Rd , Rs , Re = rates of Debt(after tax) , Retained earnings , New common stock

WACC = 0.3*5.4 + 0.1149*15 + 0.5851*18

= 1.62 + 1.7235 + 10.5318

= 13.88%(rounded to two decimals)